OpenStudy (anonymous):

How do i do this??? Let R be the region in the xy-plane between the graphs of y = ex and y = e-x from x = 0 to x = 2. a) Find the volume of the solid generated when R is revolved about the x-axis.

OpenStudy (anonymous):

did you graph the functions?

OpenStudy (anonymous):

no, what exactly do i graph?

OpenStudy (anonymous):

y=e^x , y=e^-x, and x=2

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

then you're taking the region and rotate it over the x axis - try to get the 3d shape

OpenStudy (anonymous):

you'll have to use the washer method

OpenStudy (anonymous):

Í (πR^2 - πr^2) dx , from 0 to 2

OpenStudy (anonymous):

ok thank you I see, then what?

OpenStudy (anonymous):

what is R and what is r?

OpenStudy (anonymous):

so R and r would be from top to bottom or left to right?

OpenStudy (anonymous):

imagine a quarter with a whole in the middle (washer) positioned vertically so you can stack all these washers to approximate the volume.... R is from the x axis to y=e^x, and r is from the x axis to y = e^-x

OpenStudy (anonymous):